# A Large-scale Empirical Study on Large Language Models for Election Prediction

### v1

@@ 5 Conclusion and Future Directions
Limitations
First, our approach to time-dependent modeling does not account for dynamic factors such as shifts in public opinion, evolving media narratives, or unforeseen events—such as economic disruptions or political scandals—that could significantly influence voter behavior. Incorporating these variables would require real-time data integration, which is beyond the scope of this study.
Second, while our analysis reveals potential political biases within the LLM’s pretrained corpus, we have not quantitatively measured or compared the intrinsic political tendencies of different LLMs. Without a systematic framework to quantify and contrast these inclinations, it remains challenging to understand how model-specific political leanings might influence predictions and outcomes.
It is key to emphasize that this study does not seek to provide definitive election forecasts. Rather, it serves as an exploratory effort to demonstrate the potential of LLMs to generalize to unseen data. By evaluating how closely these models align with real-world patterns, this work offers a proof of concept for their predictive capacity while acknowledging the limitations inherent in the current framework.
@@
Appendix A Extended Related Work
A.1 LLMs in Political Science: A New and Emerging Field
The application of LLMs in political science represents a new and rapidly evolving field, with a limited but growing body of research. While LLMs have revolutionized natural language processing and various other domains, their potential in political science remains largely untapped. Initial studies have demonstrated promising results in areas such as election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024). However, political science often involves complex social dynamics and multi-layered causal relationships, posing significant challenges for effectively utilizing LLMs in this context Brown (2023).
Recent work by Chen and Rodriguez (2024) highlights the potential of LLMs in decoding political speeches and policy documents, emphasizing the need for domain-specific fine-tuning to capture the subtleties of political language. Future research is likely to focus on designing models that can handle the intricacies of political discourse while ensuring robustness against biases and misleading inferences Wilson et al. (2024); Thompson (2023).
The latest work by Li et al. (2024a) presents the first systematic framework for integrating LLMs into computational political science. The authors propose a novel taxonomy that categorizes existing work into two main perspectives: political science applications and computational methodologies. From the political science perspective, they highlight LLMs’ capabilities in automating predictive and generative tasks, simulating behavior dynamics, and enhancing causal inference. From the computational perspective, they detail advances in data preparation, fine-tuning approaches, and evaluation methods specifically tailored for political contexts. The paper identifies critical challenges, including the need for domain-specific datasets, addressing bias and fairness issues, incorporating human expertise effectively, and developing evaluation criteria that are aligned with political science requirements. Our work differs from Political-LLM by giving more in-depth analysis on the election forecasting.
A.2 Political Election Research: Classical and Agent-Based Approaches
Traditional political science literature has long relied on survey data and statistical models to analyze voter behavior and predict election outcomes. Classical models like the Downsian spatial model and the Median Voter Theorem explain election results by assuming voters’ positions in policy space Downs (1957); Black (1948). Time-series regression models also play a key role in long-term election forecasting, analyzing the relationships between economic indicators, political events, and voter sentiment Lewis-Beck and Rice (1992); Erikson and Wlezien (2016).
In recent years, Agent-Based Models (ABMs) have gained traction in election research. These bottom-up approaches simulate individual voter decisions and social interactions, providing a more granular view of electoral dynamics. For instance, Gao et al. (2022a) introduced an agent-based election prediction model that captures social networks and voter-candidate interactions to offer more accurate election forecasts. This approach builds upon earlier work by Lemos et al. (2019), who demonstrated the effectiveness of ABMs in modeling voter turnout and preference formation.
ABMs excel at modeling heterogeneity and dynamic processes in real-world voting scenarios, offering greater flexibility and granularity than traditional statistical models. The work of Collins and Martinez (2023) further illustrates how ABMs can incorporate complex factors such as media influence and peer effects in voter decision-making processes. As the field evolves, future election studies may integrate ABMs with LLMs to balance large-scale data analysis with individual voter behavior modeling, as proposed by Zhang et al. (2024b) in their hybrid forecasting framework.
A.3 LLMs in Political Science: Current Research and Our Contributions
Although LLMs have seen rapid advancements, their application in political science remains limited. Only a small number of studies have explored how LLMs can be used for tasks like election prediction, policy analysis, and public opinion tracking. Political text analysis is an important area, and some early benchmark datasets are starting to emerge. However, political language is often complex, with nuanced meanings and context, which presents a significant challenge for LLMs Anderson and Taylor (2023); Williams et al. (2024).
One notable example is the “Political Campus” project by Roberts et al. (2023), which developed a benchmark dataset specifically for election prediction and evaluated LLM performance on various election-related questions. This work has been instrumental in highlighting both the potential and limitations of LLMs in political forecasting. Similarly, the research by Kim and Patel (2024) on using LLMs for policy sentiment analysis demonstrates the models’ capacity to process large volumes of public opinion data, while also underscoring the need for careful interpretation of results.
In contrast to the existing work, our research provides a more comprehensive approach. In 2024, we introduced an integrated framework that combines LLMs with Agent-Based Models to predict election outcomes and analyze voter behavior. Unlike previous studies, our approach not only focuses on semantic analysis of political texts but also incorporates individual voter behavior modeling, offering more granular and accurate election predictions. This integration addresses the limitations identified by Lopez and Singh (2023) regarding the disconnect between macro-level language models and micro-level voter behavior.
Appendix B Details of Multi-Step LLM Pipeline
B.1 Dataset Details
@@
B.1.1 Real-world Data by American National Election Studies (ANES)
@@
For evaluation, we use pre-election data from the ANES 2016 and 2020 Time Series Studies Argyle et al. (2022); Studies (2022), which provide 4,270 and 8,280 real-world samples, respectively, from individuals who participated in the 2016 and 2020 elections. The dataset includes a wide range of variables: (1) racial/ethnic self-identification, (2) gender, (3) age, (4) ideological self-placement on a conservative-liberal scale, (5) party identification, (6) political interest, (7) church attendance, (8) frequency of discussing politics with family and friends, (9) patriotic feelings associated with the American flag (unavailable in 2020), and (10) state of residence (unavailable in 2020). Additionally, the dataset records how individuals voted in both the 2016 and 2020 elections. Previous studies, such as Argyle et al. (2023), have evaluated GPT-3 using this dataset. We apply our method directly to this established benchmark to assess its effectiveness and performance.
B.1.2 Synthetic Data for the US Population
@@
In addition to the medium-sized benchmark dataset, we utilize synthetic demographic data derived from a 1:1 synthetic population dataset of the United States Li et al. (2020b). Synthetic data plays a crucial role in social and applied sciences, with recent applications in water quality estimation Chia et al. (2023), financial modeling Potluru et al. (2023a), tourist profiling Merinov et al. (2023a), and measuring the social impact of engineered products Stevenson et al. (2023). High-quality synthetic datasets provide researchers with large-scale data at a lower cost while maintaining privacy, making them a reliable resource.
@@
Synthetic Data Generation: The synthetic data used here is generated using the SynC framework Li et al. (2020b), which reconstructs individual-level data from aggregated sources where collecting real-world individual data is impractical due to privacy, time, or financial constraints. SynC is widely recognized and applied across multiple fields to support research and overcome data limitations. For instance, it has been used in outlier detection Li et al. (2020a), finance Potluru et al. (2023b), tabular data modeling Borisov et al. (2022), healthcare Sichani et al. (2024), and tourism Merinov et al. (2023b), demonstrating its effectiveness and importance in various domains.
@@
B.2 Detailed Evaluation Metrics
@@ 2. Weighted Absolute Error (WAE)
  • $P(s)$: The predicted proportion, calculated as the ratio of Republican votes to total votes (Republican + Democrat) for each state (A3).
@@ 3. Weighted Mean Squared Error (WMSE)
  • $(P(s)-R(s))^{2}$: The squared error between the predicted and actual proportions for each state.
@@ 4. Bias Metric (BM)
For the synthetic dataset, we treat each state as an independent validation unit. The predicted results—both in terms of the winning candidate and vote share percentages—are compared against the actual 2020 election results for each state. Accuracy is evaluated based on:
@@
This state-level evaluation leverages voter-level information processed through LLMs to generate accurate predictions, providing a robust assessment of model performance across diverse electoral scenarios.
Appendix C Evaluations on Synthetic Data for the 2020 US Population
Figure A1: LLM’s predictions for four states in the 2020 election compared with Ground Truth results. The figure presents results for one red state (Ohio, OH), one blue state (Illinois, IL), one swing state (Wisconsin, WI), and one tipping-point state (Florida, FL). V1 and V2 pipelines tend to underestimate Republican support, while V3 (Multi-step Reasoning) provides the closest alignment with actual outcomes, especially in swing and tipping-point states.
In addition to the nationwide evaluation on the ANES datasets, we conducted state-level simulations using synthetic data to compare predictions with actual 2020 election outcomes. For each state, we performed random sampling based on population size to ensure a statistically meaningful number of personas. The simulation outcomes were then benchmarked against official 2020 Presidential General Election Results from the Federal Election Commission (FEC). As in the benchmark evaluations, we calculated the average voting probabilities to assess the alignment of predictions with real-world outcomes. We evaluated five red states, five blue states, and 11 swing and tipping-point states. Figure A1 highlights representative results from these categories, providing insights into the model’s performance in different electoral contexts.
@@ Appendix C Evaluations on Synthetic Data for the 2020 US Population
For swing and tipping-point states, the V3 pipeline achieved robust results, correctly predicting the outcomes in 9 out of 11 states. Minor deviations were observed in North Carolina (NC) and Arizona (AZ), where the predictions were slightly misaligned with the real results. Nonetheless, the V3 pipeline provided balanced predictions that accurately captured the competitive dynamics typical of swing states, further validating its effectiveness.
In summary, the comparative performance of the three pipelines across different state categories is shown in Figure A1. The V3 pipeline consistently outperformed the other two, delivering more stable and accurate predictions. Aggregate results for all pipelines on all 21 chosen states is shown in the below figure  A2. And all state-level simulation results can be found in the Figure A3.
Appendix D Additional Results on 2024 Prediction
The 2024 state-level predictions offer deeper insights into the performance of the proposed pipelines across diverse electoral contexts. As discussed in §3.1, simulations for the 2024 election indicate a systematic bias toward the Democratic Party across the 11 swing and tipping-point states. This bias may reflect the LLM’s sensitivity to candidate-specific factors in the 2024 context. Specifically, prior to the election, V3 (§2.2.3) was evaluated across all 50 states, while V1 (§2.2.1) and V2 (§2.2.2) were tested on selected swing states and traditional red and blue states. The comparative performance of these pipelines is presented in Figure A4.
@@ Appendix D Additional Results on 2024 Prediction
Figure A3: Overall performance of the three pipelines (V1, V2, V3) in the 2020 election across all selected states. These include five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the official 2020 election results published by the Federal Election Commission Federal Election Commission (2021), while the black reference line represents an equal vote share (0.5) between the two parties.
Figure A4: Overall performance of the three pipelines (V1, V2, V3) in the 2024 election across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the 2024 election results reported by the NBC News NBC News (2024), while the black reference line represents an equal vote share (0.5) between the two parties.

### v3

@@ 5 Conclusion
Appendix A Details of Multi-Step LLM Pipeline
A.1 Dataset Details
@@
A.1.1 Real-world Data by American National Election Studies (ANES)
@@
For evaluation, we use data from the ANES 2016 and 2020 Time Series Studies Studies (2019, 2022), which provide 4,270 and 8,280 real-world samples, respectively, from individuals who participated in the 2016 and 2020 elections. The dataset includes a wide range of variables: (1) racial/ethnic self-identification, (2) gender, (3) age, (4) ideological self-placement on a conservative-liberal scale, (5) party identification, (6) political interest, (7) church attendance, (8) frequency of discussing politics with family and friends, (9) patriotic feelings associated with the American flag (unavailable in 2020), and (10) state of residence (unavailable in 2020). Additionally, the dataset records how individuals voted in both the 2016 and 2020 elections. Previous studies, such as Argyle et al. (2023), have evaluated GPT-3 using this dataset. We apply our method directly to this established benchmark to assess its effectiveness and performance.
A.1.2 Synthetic Personas for the U.S. Population
@@
In addition to the medium-sized benchmark dataset, we utilize synthetic demographic data derived from a 1:1 synthetic population dataset of the United States Li et al. (2020b). Synthetic data plays a crucial role in social and applied sciences, with recent applications in water quality estimation Chia et al. (2023), financial modeling Potluru et al. (2023a), tourist profiling Merinov et al. (2023), and measuring the social impact of engineered products Stevenson et al. (2023). High-quality synthetic datasets provide researchers with large-scale data at a lower cost while maintaining privacy, making them a reliable resource.
@@
Synthetic Data Generation: The synthetic data used here is generated using the SynC framework Li et al. (2020b), which reconstructs individual-level data from aggregated sources where collecting real-world individual data is impractical due to privacy, time, or financial constraints. SynC is widely recognized and applied across multiple fields to support research and overcome data limitations. For instance, it has been used in outlier detection Li et al. (2020a), finance Potluru et al. (2023b), tabular data modeling Borisov et al. (2022), healthcare Sichani et al. (2024), and tourism Merinov et al. (2023), demonstrating its effectiveness and importance in various domains.
@@
A.2 Detailed Evaluation Metrics
@@ 2. Weighted Absolute Error (WAE)
  • $P(s)$: The simulated proportion, calculated as the ratio of Republican votes to total votes (Republican + Democrat) for each state (A3).
@@ 3. Weighted Mean Squared Error (WMSE)
  • $(P(s)-R(s))^{2}$: The squared error between the simulated and actual proportions for each state.
@@ 4. Bias Metric (BM)
For the synthetic dataset, we treat each state as an independent validation unit. The simulated results—both in terms of the winning candidate and vote share percentages—are compared against the actual 2020 election results for each state. Accuracy is evaluated based on:
@@
This state-level evaluation leverages voter-level information processed through LLMs to generate accurate simulations, providing a robust assessment of model performance across diverse electoral scenarios.
Appendix B Evaluations on Synthetic Personas for the 2020 U.S. Population
Figure A1: LLM’s simulations for four states in the 2020 election compared with Ground Truth results. The figure presents results for one red state (Ohio, OH), one blue state (Illinois, IL), one swing state (Wisconsin, WI), and one tipping-point state (Florida, FL). V1 and V2 pipelines tend to underestimate Republican support, while V3 (Multi-step Reasoning) provides the closest alignment with actual outcomes, especially in swing and tipping-point states.
In addition to the nationwide evaluation on the ANES datasets, we conducted state-level simulations using synthetic data to compare it with actual 2020 election outcomes. For each state, we performed random sampling based on population size to ensure a statistically meaningful number of personas. The simulation outcomes were then benchmarked against official 2020 Presidential General Election Results from the Federal Election Commission (FEC). As in the benchmark evaluations, we calculated the average voting probabilities to assess the alignment of predictions with real-world outcomes. We evaluated five red states, five blue states, and 11 swing and tipping-point states. Figure A1 highlights representative results from these categories, providing insights into the model’s performance in different electoral contexts.
@@
For swing and tipping-point states, the V3 pipeline achieved robust results, correctly simulating the outcomes in 9 out of 11 states. Minor deviations were observed in North Carolina (NC) and Arizona (AZ), where the predictions were slightly misaligned with the real results. Nonetheless, the V3 pipeline provided balanced predictions that accurately captured the competitive dynamics typical of swing states, further validating its effectiveness.
In summary, the comparative performance of the three pipelines across different state categories is shown in Figure A1. The V3 pipeline consistently outperformed the other two, delivering more stable and accurate predictions. Aggregate results for all pipelines on all 21 chosen states is shown in the below figure  A2.
Appendix C Additional Results on 2024 Prediction
The 2024 state-level Simulations offer deeper insights into the performance of the proposed pipelines across diverse electoral contexts. As discussed in §3.1, simulations for the 2024 election indicate a systematic bias toward the Democratic Party across the 11 swing and tipping-point states. This bias may reflect the LLM’s sensitivity to candidate-specific factors in the 2024 context. Specifically, prior to the election, V3 (§2.2.3) was evaluated across all 50 states, while V1 (§2.2.1) and V2 (§2.2.2) were tested on selected swing states and traditional red and blue states. The comparative performance of these pipelines is presented in Figure A6.
@@ Appendix C Additional Results on 2024 Prediction
Appendix D Beyond Accuracy
Figure A3: Demographic (Age) voting pattern: comparing LLM Simulations with real human data (Pew Report)
Figure A4: Demographic (Education) voting pattern: comparing LLM Simulations with real human data (Pew Report)
Appendix E Broader Impact Statement
This work explores the application of LLMs to simulate voter behaviour through enhanced reasoning and data synthesis, which may inform policymakers, researchers, and journalists, aiding them in understanding voter behavior and electoral outcomes. By offering a more transparent and adaptable approach to prediction, this research may help demystify complex political processes, reduce reliance on narrow historical data, and guide strategic resource allocation for stakeholders.
Figure A5: Overall performance of the three pipelines (V1, V2, V3) in the 2020 simulation across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the 2020 election results Federal Election Commission (2021), while the black reference line represents an equal vote share (0.5) between the two parties.
Figure A6: Overall performance of the three pipelines (V1, V2, V3) in the 2024 simulation across five red states (Arkansas (AR), South Carolina (SC), Idaho (ID), Alabama (AL), Ohio (OH)), five blue states (Illinois (IL), California (CA), New York (NY), New Jersey (NJ), Washington (WA)), 11 swing and tipping-point states (New Hampshire (NH), Arizona (AZ), Pennsylvania (PA), Georgia (GA), Minnesota (MN), North Carolina (NC), Wisconsin (WI), Florida (FL), Michigan (MI), Nevada (NV), Texas (TX)), and an additional red state (Alaska (AK)). The red reference line corresponds to the 2024 election results reported by the NBC News NBC News (2024), while the black reference line represents an equal vote share (0.5) between the two parties.